How many self dual functions are possible with four boolean variables?
1
Expert's answer
2014-01-20T03:50:16-0500
Theorem: There are 2^(2^(n-1)) different self-dual functions of n variables.
Using this theorem we get
2^(2^(4-1))=2^(2^3)=2^8=256 functions.
Answer: 256 functions.